Cameron has 11 adult fantail goldfish. This is 7 fewer fantail goldfish than twice his friend Julia has. Write and solve an equation to determine the number of fantail goldfish that Julia has.
Answer by TimothyLamb(4379)

x = number of fish cameron
y = number of fish julia
---
x = 11
x = 2y - 7
---
put the system of linear equations into standard form
---
x = 11
x - 2y = -7
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = number of fish cameron = 11
y = number of fish julia = 9
---
now by substitution:
---
x = 11
x = 2y - 7
11 = 2y - 7
2y = 11 + 7
2y = 18
y = 18/2
y = 9
---
answer by substitution:
x = number of fish cameron = 11
y = number of fish julia = 9
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---